RESIDUALS IN THE GROWTH CURVE MODEL

DIETRICH VON ROSEN

Department of Mathematics, Uppsala University, Box 480, S-751 06 Uppsala, Sweden

(Received April 8, 1994; revised October 11, 1994)

Abstract. Residuals for the Growth Curve model will be discussed. In univariate linear models as well as the ordinary multivariate analysis of variance model residuals are based on the difference between the observations and the mean whereas for the Growth Curve model we have three different residuals all showing various aspects useful for validating analysis. For these residuals some basic properties are established.

Key words and phrases: GMANOVA, Growth Curve model, residuals.